Suppose Cor(X,Y)=1/3 and oy = 403. Let Z=2X +3Y. If Var(Z) = 240, calculate Var(X).
Answer #1

Var(Z) = Var(2X + 3Y) = 4*Var(X) + 9*Var(Y) + 2*2*3*Cov(X, Y)

= 4*Var(X) + 9*Var(Y) + 2*2*3*Corr(X, Y)*SD(X)*SD(Y) ----- (1)

[SD means standard deviation. Note that Corr(X, Y) = Cov(X, Y)/SD(X)SD(Y)]

Given that Var(Y) = 4*Var(X). Hence, SD(Y) = \sqrt{Var(Y)} = 2*\sqrt{Var(X)} = 2*SD(X) Replacing this in above eqution (1), we get Var(Z) = 4*Var(X) + 9*4*Var(X) + 2*2*3*(1/3)*SD(X)*2*SD(X)

Or, 240 = 40*Var(X) + 8*SD(X)*SD(X) = 40*Var(X) + 8*Var(X) = 48*Var(X)

Hence, Var(X) = 240/48 = 5